OpenStudy (anonymous):

CosA+CosB-CosC = -1+4CosA/2. CosB/2.SinC/2

OpenStudy (thomas5267):

\[ \cos(A)+\cos(B)-\cos(C)=-1+4\cos\left(\frac{A}{2}\right)\cos\left(\frac{B}{2}\right)\sin\left(\frac{C}{2}\right) \] Is that the equation?

OpenStudy (anonymous):

A+B+C=180 A=180-(B+C) cos A=cos{180-(B+C)}=-cos(B+C)=\[-[1-2\sin ^2\frac{ B+C }{ 2 }]\] \[=-1+2\sin ^2\frac{ B+C }{ 2 }\] \[\cos2x=1-2\sin ^2x\] \[\cos x-\cos y=-2\sin \frac{ x+y }{ 2 }\sin \frac{ x-y }{ 2 }\] use it and complete it.